The manager of this retail stores decides to statistically analyse the relationship between sales of umbrellas (SALES) and the daily maximum temperature (MAXTEMP) with the number of customers (CUSTOMERS) visiting the store, irrespective of whether the customer purchased an umbrella.The regression model is: SALES(predicted)=472 - 4*MAXTEMP +3.5*CUSTOMERS. What is the expected sales on a day when MAXTEMP = 19C and CUSTOMERS = 220

Select one:

a. 1166

b. 6671

c. 2352

d. 472

e. 2743
Answer #1

Given,

Expected Sales = 472 - 4*MAXTEMP +3.5*CUSTOMERS

MAXTEMP=19C

CUSTOMERS = 220

Substituting the given values in expected sales,

Expected Sales = 472 -4*19 + 3.5 * 220 = 472-76+770 = 1166

Hence, the correct answer is a. 1166